# Why Policy Gradient Algorithms Work for Undiscounted Total-Reward MDPs

### v1

@@ 1 Introduction
Contribution
In this work, we study the convergence of policy gradient methods for undiscounted expected total-reward infinite-horizon MDPs. Our analysis is based on two key insights: (i) the classification of the MDP states into recurrent and transient states is invariant over the set of policies that assign strictly positive probability to every action (as is typical in deep RL models employing a softmax output layer) and (ii) the classical state visitation measure (which may ill-defined when $\gamma=1$) can be replaced with a new object that we call the transient visitation measure. Leveraging these insights, we establish convergence guarantees for projected policy gradient and natural policy gradient algorithms in the tabular setting.
@@
Undiscounted total-reward infinite horizon MDP.
The setup of undiscounted total-reward MDP was first introduced by Savage (1965). For the well-definedness of the value function, Schäl (1983) considered the finiteness of $V^{\pi}_{+}$ and $V^{\pi}_{-}$ (defined in the next section) and the existence of an optimal policy was first proved by Van Der Wal (1981). With different additional assumptions, three models of total-reward setup have been proposed and studied: stochastic shortest path model (Eaton & Zadeh, 1962), positive model (Blackwell, 1967), and negative model (Strauch, 1966). The stochastic shortest path model assumes single absorbing terminal state and the existence of a policy that reaches the terminal state with probability $1$ from any initial state (Bertsekas & Tsitsiklis, 1991). The positive model assumed $V_{+}^{\pi}$ is finite, and for each $s$, there exist $a$ with $r(s,a)\geq 0$ (Puterman, 2014, Section 7.2). The negative model assumed $V_{+}^{\pi}=0$ and there exist $\pi$ for which $V_{-}^{\pi}(s)>\infty$ for all $s$ (Puterman, 2014, Section 7.3). In this work, the undiscounted total-reward model with the assumption that $V^{\pi}$ is finite, motivated by modern reinforcement learning frameworks and needed to establish the transient policy gradient, does not fall into previous categories.
@@
Policy gradient methods (Williams, 1992; Sutton et al., 1999; Konda & Tsitsiklis, 1999; Kakade, 2001) are foundational reinforcement learning algorithms, commonly implemented with deep neural networks for policy parameterization (Schulman et al., 2015; 2017). In line with their practical success, convergence of policy gradient variants has been extensively studied across settings. In discounted total-reward infinite horizon MDP, Agarwal et al. (2021); Xiao (2022); Bhandari & Russo (2024); Mei et al. (2020) analyzed convergence of projected policy gradient and naive policy gradient with softmax parametrization. The natural policy gradient, introduced by Kakade (2001) and viewable as a special case of mirror descent (Shani et al., 2020), has been analyzed by Agarwal et al. (2021); Cen et al. (2022); Xiao (2022); Lan (2023). In the average reward MDP, convergence results have been established by Even-Dar et al. (2009); Murthy & Srikant (2023); Bai et al. (2024); Kumar et al. (2024), and related analyses exist for the finite horizon setup as well (Hambly et al., 2021; Guo et al., 2022; Klein et al., 2023).
In undiscounted total-reward infinite-horizon MDP, however, there are few results on policy gradient methods. Since Sutton et al. (1999) established the policy gradient theorem only for the discounted total-reward and average reward MDPs, Bojun (2020); Ribera Borrell et al. (2025) analyze policy gradients for the undiscounted total-reward random time horizon MDP. Specifically, Bojun (2020) consider an episodic learning process that can be viewed as an ergodic Markov chain with finite episode length , and establish a policy gradient theorem via the steady state distribution. Ribera Borrell et al. (2025) consider trajectory dependent random termination times and prove a policy gradient theorem with an almost surely finite termination time. We note that neither work further analyzes the convergence of policy gradient methods, and their setups and assumptions differ from ours, as shown in next section.
@@ 2 Undiscounted expected total-reward infinite-horizon MDPs
In this work, we consider undiscounted total-reward infinite-horizon Markov decision processes (MDPs). We review basic definitions and assumptions of undiscounted MDPs and reinforcement learning (RL). For further details, we refer the readers to references such as (Puterman, 2014, Section 7) or (Sutton & Barto, 2018).
@@
Assumption 1 (Finiteness of value function).
@@
As noted, in recent training of large language models under the frameworks of reinforcement learning with human feedback (RLHF) and reinforcement learning with verifiable reward (RLVR), finite rewards are assigned once at the end of the trajectory, and Assumption 1 holds.
Under Assumption 1, we define the state value function as
@@
and $Q_{\raisebox{1.35623pt}{$\mkern 1.0mu\scriptscriptstyle-$}}^{\pi}$ analogously. Likewise, define $Q^{\pi}=Q_{\raisebox{1.35623pt}{$\mkern 1.0mu\scriptscriptstyle+$}}^{\pi}-Q_{\raisebox{1.35623pt}{$\mkern 1.0mu\scriptscriptstyle-$}}^{\pi}$. Then, $Q^{\pi}$ is well-defined under Assumption 1, and $Q^{\pi}=PV^{\pi}+r$, and $V^{\pi}(s)=\mathbb{E}_{a\sim\pi(\cdot\,|\,s)}[Q^{\pi}(s,a)]$ by definition.
We say $V^{\star}$ is optimal value function if $V^{\star}(s)=\max_{\pi}V^{\pi}(s)$ for all $s\in\mathcal{S}$ and $\pi$ is an $\epsilon$-optimal policy if $\left\|{V^{\star}-V^{\pi}}\right\|_{\infty}\leq\epsilon$. It is known that the optimal value function and an optimal policy always exist in the undiscounted total-reward setup with finite state and action spaces (Puterman, 2014, Theorem 7.19). (As a technical detail, Theorem 7.19 of Puterman (2014) is stated in terms of the set of all history-dependent policies, but the proof also works for the set of all stationary policies, which is our focus in this work.)
For notational conciseness, we write $r^{\pi}(s)=\mathbb{E}_{a\sim\pi(\cdot\,|\,s)}\left[r(s,a)\right]$ for the reward induced by policy $\pi$ and $P^{\pi}(s)$ defined as
$P^{\pi}(s\rightarrow s^{\prime})=\mathrm{Pr}(s\rightarrow s^{\prime}\,|\,a\sim\pi(\cdot\,|\,s),s^{\prime}\sim P(\cdot\,|\,s,a))$
@@ 3 Troubles with undiscounted total-reward MDPs
The value function $V^{\pi}$ may be a discontinuous function of $\pi$, even when $|\mathcal{S}|$ and $|\mathcal{A}|$ are finite and $V^{\pi}$ is finite.
In the example of Figure 1, the optimal action at state $s_{1}$ is to remain at $s_{1}$. Under the optimal policy, taking this optimal action, we have $V^{\star}(s_{1})=0$, but any policy assigning a non-zero probability to the other action, transitioning to $s_{3}$, yields $V^{\pi}(s_{1})=-1$. This example illustrates that the value function can be discontinuous in $\pi$, and a policy gradient method cannot be expected to succeed in the presence of such discontinuities. We provide the framework to address this pathology in Section 4.
@@ 4.1 Recurrent-transient classification of states
where $\bar{R}^{\pi}$, $\bar{T}^{\pi}$ represent transition probabilities between the recurrent states and the transient states, respectively. This canonical representation exists for any $\pi\in\Pi$, but the recurrent-transient classification of states and the corresponding canonical decomposition of $P^{\pi}$ may vary.
@@
The canonical recurrent-transient decomposition provides the foundation for our analysis of the undiscounted expected total-reward setting. One key consequence of this classification is that the reward-to-go at any recurrent state must be zero if the value functions are finite.
Lemma 1.
Under Assumption 1 (finiteness of value function), for any $\pi\in\Pi$, if $s$ is a recurrent state, then $r^{\pi}(s)=0$.
@@
$T^{\pi}=\begin{bmatrix}0&0\\ 0&\bar{T}^{\pi}\end{bmatrix},\qquad\text{i.e.,}\qquad T^{\pi}(s_{1},s_{2})=\left\{\begin{array}[]{ll}P^{\pi}(s_{1},s_{2})&\text{if $s_{1},s_{2}$ are both transient}\\ 0&\text{otherwise.}\end{array}\right$.
@@
(Berman & Plemmons, 1994, Lemma 8.3.20) Spectral radius of $T^{\pi}$ is strictly less than $1$.
@@
By Lemma 1, we have $(P^{\pi})^{i}r^{\pi}=(T^{\pi})^{i}r^{\pi}$ for $i\in\mathbb{N}$, which implies
@@
By Fact 1 and the classical Neumann series argument, we have $(I-T^{\pi})^{-1}=\sum_{i=0}^{\infty}(T^{\pi})^{i}$. These lead the following reformulation of the value function.
Lemma 2.
Under Assumption 1, $V^{\pi}=(I-T^{\pi})^{-1}r^{\pi}$ for any $\pi\in\Pi$.
@@ 4.2 Continuity of $V^{\pi}$ on $\Pi_{+}$
Lemma 3.
Under Assumption 1, the mappings $\pi\mapsto V^{\pi}$ and $\pi\mapsto V^{\pi}_{\mu}$ are continuous on $\Pi_{+}$ for a given $\mu$.
@@
$V^{\star}_{+}=\sup_{\pi\in\Pi_{+}}V^{\pi}\quad\text{and}\quad V_{+,\mu}^{\star}=\mathbb{E}_{s\sim\mu}[V_{+}^{\star}(s)],$
while recalling
$V^{\star}=\max_{\pi\in\Pi}V^{\pi}\quad\text{and}\quad V_{\mu}^{\star}=\mathbb{E}_{s\sim\mu}[V^{\star}(s)].$
By definition, $V^{\star}(s)\geq V^{\star}_{+}(s)$ for all $s\in\mathcal{S}$ and $V_{\mu}^{\star}\geq V_{+,\mu}^{\star}$ for any initial distribution $\mu$. Since our policy gradient methods search over $\Pi_{+}$, they should be thought of as optimizing for $V_{+,\mu}^{\star}$. However, the distinction between $V_{+,\mu}^{\star}$ and $V_{\mu}^{\star}$ disappears when rewards are nonnegative.
Lemma 4.
Under Assumption 1, if rewards are nonnegative, then the map $\pi\mapsto V^{\pi}_{\mu}$ are continuous at optimal policies, and $V^{\star}_{+,\mu}=V^{\star}_{\mu}$ for a given $\mu$.
@@ 4.3 Transient visitation measure
Conventionally, the state visitation measure is defined as $d_{s_{0}}^{\pi}(s)=(1-\gamma)\sum_{i=0}^{\infty}(\gamma P^{\pi})^{i}(s_{0},s_{i})$ in the discounted infinite-horizon setting with $\gamma<1$. In the undiscounted setting with $\gamma=1$, this object becomes undefined.
@@
$\delta^{\pi}_{s_{0}}(s)=\sum_{i=0}^{\infty}\Pr(s_{i}=s\mid s_{0};\,T^{\pi})=e_{s_{0}}^{\intercal}(I-T^{\pi})^{-1}e_{s},$
where $e_{s}$ is the $s$-coordinate vector. Also define $\delta^{\pi}_{\mu}=\mathbb{E}_{s_{0}\sim\mu}[\delta^{\pi}_{s_{0}}]$ for an initial state distribution $\mu$. Note that this transient visitation measure is not a probability measure, and, importantly, $\max_{s,s_{0}\in\mathcal{S}}\delta^{\pi}_{s_{0}}(s)<\infty$ by Fact 1. The transient visitation measure only considers transitions between transient states since these are sufficient to compute the value function, as shown in Lemma 2.
@@
Lemma 5 (Transient performance difference lemma).
Under Assumption 1, for $\pi,\pi^{\prime}\in\Pi$ and a given $\mu$, if $V^{\pi^{\prime}}(s)=0$ for all recurrent states $s$ of $P^{\pi}$, then
@@ 4.4 Transient policy gradient
Theorem 1 (Transient policy gradient).
Under Assumption 1, for $\pi_{\theta}\in\Pi_{+}$,
$\nabla_{\theta}V^{\pi_{\theta}}_{\mu}=\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi_{\theta}}(s)\sum_{a\in\mathcal{A}}\nabla_{\theta}\pi_{\theta}(a\mid s)Q^{\pi_{\theta}}(s,a)=\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi_{\theta}}(s)\mathop{\mathbb{E}}_{a\sim\pi(\cdot\,|\,s)}[\nabla_{\theta}\log\pi_{\theta}(a\mid s)Q^{\pi_{\theta}}(s,a)].$

### v2

@@ 1 Introduction
Contribution.
In this work, we study the convergence of policy gradient methods for undiscounted expected total-reward infinite-horizon MDPs. Our analysis is based on two key insights: (i) the classification of the MDP states into recurrent and transient states is invariant over the set of policies that assign strictly positive probability to every action (as is typical in deep RL models employing a softmax output layer) and (ii) the classical state visitation measure (which may be ill-defined when $\gamma=1$) can be replaced with a new object that we call the transient visitation measure. Leveraging these insights, we establish convergence guarantees for projected policy gradient and natural policy gradient algorithms in the tabular setting. We also extend the analysis to the generative model setting and derive sample complexity for stochastic natural policy gradient.
@@ 1.1 Related works
Undiscounted total-reward infinite-horizon MDP.
The setup of undiscounted total-reward MDP was first introduced by (Savage, 1965). For the well-definedness of the value function, (Schäl, 1983) considered the finiteness of $V^{\pi}_{+}$ and $V^{\pi}_{-}$ (defined in the next section) and the existence of an optimal policy was first proved by (Van Der Wal, 1981). With different additional assumptions, three models of total-reward setup have been proposed and studied: stochastic shortest path model (Eaton and Zadeh, 1962), positive model (Blackwell, 1967), and negative model (Strauch, 1966). The stochastic shortest path model assumes single absorbing terminal state and the existence of a policy that reaches the terminal state with probability $1$ from any initial state (Bertsekas and Tsitsiklis, 1991). The positive model assumed $V_{+}^{\pi}$ is finite, and for each $s$, there exist $a$ with $r(s,a)\geq 0$ (Puterman, 2014, Section 7.2). The negative model assumed $V_{+}^{\pi}=0$ and there exist $\pi$ for which $V_{-}^{\pi}(s)>\infty$ for all $s$ (Puterman, 2014, Section 7.3). In this work, the undiscounted total-reward model with the assumption that $V^{\pi}$ is finite, motivated by modern reinforcement learning frameworks and needed to establish the transient policy gradient, does not fall into previous categories.
@@
Policy gradient methods (Williams, 1992; Sutton et al., 1999; Konda and Tsitsiklis, 1999; Kakade, 2001) are foundational reinforcement learning algorithms, commonly implemented with deep neural networks for policy parameterization (Schulman et al., 2015, 2017). In line with their practical success, convergence and sample complexity of policy gradient variants have been extensively studied across settings. In discounted total-reward infinite-horizon MDP, (Agarwal et al., 2021; Xiao, 2022; Bhandari and Russo, 2024; Mei et al., 2020) analyzed convergence of projected policy gradient and naive policy gradient with softmax parametrization. The natural policy gradient, introduced by (Kakade, 2001) and viewable as a special case of mirror descent (Shani et al., 2020), has been analyzed by (Agarwal et al., 2021; Cen et al., 2022; Xiao, 2022; Lan, 2023). In the average reward MDP, where the goal is to maximize long-term, steady-state performance, convergence and sample complexity results have been established by (Even-Dar et al., 2009; Murthy and Srikant, 2023; Bai et al., 2024; Kumar et al., 2024; Li et al., 2025), and related analyses exist for the finite horizon setup as well (Hambly et al., 2021; Guo et al., 2022; Klein et al., 2023).
In undiscounted total-reward infinite-horizon MDP, however, there are few results on policy gradient methods. Since Sutton et al. (1999) established the policy gradient theorem only for the discounted total-reward and average reward MDPs, Bojun (2020); Ribera Borrell et al. (2025) analyze policy gradients for the undiscounted total-reward random time horizon MDP. Specifically, Bojun (2020) considers an episodic learning process that can be viewed as an ergodic Markov chain with finite episode length, and establish a policy gradient theorem via the steady state distribution. Ribera Borrell et al. (2025) consider trajectory dependent random termination times and prove a policy gradient theorem with an almost surely finite termination time. We note that neither work further analyzes the convergence of policy gradient methods, and their setups and assumptions differ from ours, as shown in the next section.
@@ 2 Undiscounted expected total-reward infinite-horizon MDPs
In this work, we consider undiscounted total-reward infinite-horizon Markov decision processes (MDPs). We review basic definitions and assumptions of undiscounted MDPs and reinforcement learning (RL). For further details, we refer the readers to references such as (Puterman, 2014, Section 7) or (Sutton and Barto, 2018).
@@
Assumption 2.1 (Finiteness of value function).
@@
While bounded rewards ensure the well-definedness of the value function and policy gradient in discounted and average-reward settings, our undiscounted total-reward setup requires Assumption 2.1. As noted, in recent training of large language models under the frameworks of reinforcement learning with human feedback (RLHF) and reinforcement learning with verifiable reward (RLVR), finite rewards are assigned once at the end of the trajectory, and Assumption 2.1 naturally holds.
Under Assumption 2.1, we define the state value function as
@@
and $Q_{\raisebox{1.35623pt}{$\mkern 1.0mu\scriptscriptstyle-$}}^{\pi}$ analogously. Likewise, define $Q^{\pi}=Q_{\raisebox{1.35623pt}{$\mkern 1.0mu\scriptscriptstyle+$}}^{\pi}-Q_{\raisebox{1.35623pt}{$\mkern 1.0mu\scriptscriptstyle-$}}^{\pi}$. Then, $Q^{\pi}$ is well-defined under Assumption 2.1, and $Q^{\pi}=PV^{\pi}+r$ where $P\in^{|\mathcal{S}|\!|\mathcal{A}|\!\times\!|\mathcal{S}|}$, and $V^{\pi}(s)=\mathbb{E}_{a\sim\pi(\cdot\,|\,s)}[Q^{\pi}(s,a)]$ by definition.
We say $V^{\star}$ is optimal value function if $V^{\star}(s)=\max_{\pi}V^{\pi}(s)$ for all $s\in\mathcal{S}$ and $\pi$ is an $\epsilon$-optimal policy if $\left\|{V^{\star}-V^{\pi}}\right\|_{\infty}\leq\epsilon$. It is known that the optimal value function and an optimal policy always exist in the undiscounted total-reward setup with finite state and action spaces (Puterman, 2014, Theorem 7.1.9). (As a technical detail, Theorem 7.1.9 of (Puterman, 2014) is stated in terms of the set of all history-dependent policies, but the proof also works for the set of all stationary policies, which is our focus in this work.)
For notational conciseness, we write $r^{\pi}(s)=\mathbb{E}_{a\sim\pi(\cdot\,|\,s)}\left[r(s,a)\right]$ for the reward induced by policy $\pi$ and $P^{\pi}(s,s^{\prime})$ defined as
$P^{\pi}(s,s^{\prime})=\mathrm{Prob}(s\rightarrow s^{\prime}\,|\,a\sim\pi(\cdot\,|\,s),s^{\prime}\sim P(\cdot\,|\,s,a))$
@@ 3 Troubles with undiscounted total-reward MDPs
The value function $V^{\pi}$ may be a discontinuous function of $\pi\mathrel{\in}\Pi$, even when $|\mathcal{S}|$ and $|\mathcal{A}|$ are finite and $V^{\pi}$ is finite.
In the example of Figure 1, the optimal action at state $s_{1}$ is to remain at $s_{1}$. Under the optimal policy, taking this optimal action, we have $V^{\star}(s_{1})=0$, but any policy assigning a non-zero probability to the other action, transitioning to $s_{3}$, yields $V^{\pi}(s_{1})=-1$. This example illustrates that the value function can be discontinuous in $\pi$, and a policy gradient method cannot be expected to succeed in the presence of such discontinuities, while the discounted-reward setup guarantees differentiability of the value function. We provide the framework to address this pathology in Section 4.
@@ 4.1 Recurrent-transient classification of states
where $\bar{R}^{\pi}$, $\bar{T}^{\pi}$, and $\bar{S}^{\pi}$ represent transition probabilities among the recurrent states, among the transient states, and from transient to recurrent states, respectively. This canonical representation exists for any $\pi\in\Pi$, but the recurrent-transient classification of states and the corresponding canonical decomposition of $P^{\pi}$ may vary.
@@
The canonical recurrent-transient decomposition provides the foundation for our analysis of the undiscounted expected total-reward setting. One key consequence of this classification is that the reward at any recurrent state must be zero if the value functions are finite.
Lemma 4.2.
Under Assumption 2.1 (finiteness of value function), for any $\pi\in\Pi$, if $s$ is a recurrent state, then $r^{\pi}(s)=0$.
@@
$T^{\pi}=\begin{bmatrix}0&0\\ 0&\bar{T}^{\pi}\end{bmatrix}$
i.e.,
$T^{\pi}(s_{1},s_{2})=\left\{\begin{array}[]{ll}P^{\pi}(s_{1},s_{2})&\text{if $s_{1},s_{2}$ are both transient}\\ 0&\text{otherwise.}\end{array}\right$.
@@
(Berman and Plemmons, 1994, Lemma 8.3.20) Spectral radius of $T^{\pi}$ is strictly less than $1$.
@@
By Lemma 4.2, we have $(P^{\pi})^{i}r^{\pi}=(T^{\pi})^{i}r^{\pi}$ for $i\in\mathbb{N}$, which implies
@@
By Fact 1 and the classical Neumann series argument, we have $(I-T^{\pi})^{-1}=\sum_{i=0}^{\infty}(T^{\pi})^{i}$. These lead to the following reformulation of the value function.
Lemma 4.3.
Under Assumption 2.1, $V^{\pi}=(I-T^{\pi})^{-1}r^{\pi}$ for any $\pi\in\Pi$.
@@ 4.2 Continuity of $V^{\pi}$ on $\Pi_{+}$
Lemma 4.4.
Under Assumption 2.1, the mappings $\pi\mapsto V^{\pi}$ and $\pi\mapsto V^{\pi}_{\mu}$ are continuous on $\Pi_{+}$ for a given $\mu$.
@@
$V_{+,\mu}^{\star}=\sup_{\pi\in\Pi_{+}}\mathbb{E}_{s\sim\mu}[V^{\pi}(s)],\qquad V_{\mu}^{\star}=\max_{\pi\in\Pi}\mathbb{E}_{s\sim\mu}[V^{\pi}(s)],$
where $\mu$ is an initial state distribution. By definition, $V_{\mu}^{\star}\geq V_{+,\mu}^{\star}$. Since our policy gradient methods search over $\Pi_{+}$, they should be thought of as optimizing for $V_{+,\mu}^{\star}$. However, the distinction between $V_{+,\mu}^{\star}$ and $V_{\mu}^{\star}$ disappears when rewards are nonnegative.
Lemma 4.5.
Under Assumption 2.1, if rewards are nonnegative, then the map $\pi\mapsto V^{\pi}_{\mu}$ are continuous at optimal policies, and $V^{\star}_{+,\mu}=V^{\star}_{\mu}$ for a given $\mu$.
@@ 4.3 Transient visitation measure
Conventionally, the state visitation measure is defined as $d_{s_{0}}^{\pi}(s)=(1-\gamma)\sum_{i=0}^{\infty}\mathrm{Prob}(s_{i}=s\mid s_{0};\,P^{\pi})$ in the discounted infinite-horizon setting with $\gamma<1$. In the undiscounted setting with $\gamma=1$, this object becomes undefined.
@@
$\delta^{\pi}_{s_{0}}(s)=\sum_{i=0}^{\infty}\mathrm{Prob}(s_{i}=s\mid s_{0};\,T^{\pi})=e_{s_{0}}^{\intercal}(I-T^{\pi})^{-1}e_{s},$
where $e_{s}$ is the $s$-coordinate vector. Also define $\delta^{\pi}_{\mu}=\mathbb{E}_{s_{0}\sim\mu}[\delta^{\pi}_{s_{0}}]$ for an initial state distribution $\mu$. Note that this transient visitation measure is not a probability measure, and, importantly, $\max_{s,s_{0}\in\mathcal{S}}\delta^{\pi}_{s_{0}}(s)<\infty$ by Fact 1. The transient visitation measure only considers transitions between transient states since these are sufficient to compute the value function, as shown in Lemma 4.3.
@@
Lemma 4.6 (Transient performance difference lemma).
Under Assumption 2.1, for $\pi,\pi^{\prime}\in\Pi$ and a given $\mu$, if $V^{\pi^{\prime}}(s)=0$ for all recurrent states $s$ of $P^{\pi}$, then
@@ 4.4 Transient policy gradient
Theorem 4.7 (Transient policy gradient).
Under Assumption 2.1, for $\pi_{\theta}\in\Pi_{+}$,
$\nabla_{\theta}V^{\pi_{\theta}}_{\mu}$ $=\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi_{\theta}}(s)\sum_{a\in\mathcal{A}}\nabla_{\theta}\pi_{\theta}(a\mid s)Q^{\pi_{\theta}}(s,a)$
$=\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi_{\theta}}(s)\mathop{\mathbb{E}}_{a\sim\pi_{\theta}(\cdot\,|\,s)}[\nabla_{\theta}\log\pi_{\theta}(a\mid s)Q^{\pi_{\theta}}(s,a)].$